# DALex: Lexicase-like Selection via Diverse Aggregation

### 23 Jan 2024 v1

@@ 3.3 Modifications
For domains suited for a relaxed lexicase variant such as epsilon-lexicase or batch-lexicase, we first standardize the population errors on each training case to have a mean of 0 and a variance of 1 before multiplying by the training case weights. While not necessary, we find that this normalization step helps DALex perform well across many problems with a single hyperparameter setting.
Of the problem domains we study, Learning Classifier Systems (LCS) [46] has a significantly different structure and therefore requires further modification. In LCS, where an individual represents a rule that matches a subset of the training cases, each individual will only have errors defined on a subset of the training cases. For these problems, we let the individual have error 0 on cases for which it is undefined. We also use the individual’s support vector, which has a 1 for each training case on which the individual is defined, and a 0 otherwise. To compute the individual’s average error, we take the dot product of the error vector with the weight vector and then normalize by dividing this value by the dot product of the support vector with the weight vector.
Algorithm 1: Diversely Aggregated Lexicase, batched selection
@@
Output:
if Relaxed:
@@
$W$leftarrow ${\tt softmax}(I,{\tt axis}=1)$ The training case weights
$F$leftarrow $\frac{EW^{T}}{SW^{T}}$ The normalized, weighted, average error for each individual
${\tt idx}$leftarrow ${\tt argmin}(F,{\tt axis=0})$ lowest error individuals w.r.t the weights $W$
return idx
In problems where individuals are defined on all training cases, the support is 1 everywhere, so the normalization constant $SW^{T}$ is simply the sum of the training case weights, which is 1 due to the softmax operation. Under this condition, our algorithm degenerates to the simpler case considered above. The complete algorithm pseudocode, with the augmentations described above to accommodate for LCS and SR problems, is described in Algorithm 1.
@@ 4.1 CBGP
We first compare DALex to lexicase selection on program synthesis problems, the class of problem for which lexicase selection was first developed. For these experiments, we use a particularity pressure of 20 As DALex can exactly simulate lexicase selection in the limit of infinite particularity pressure, we also assess the ability of DALex to replicate the lexicase selection probability distribution with a moderate particularity pressure. To situate our results with respect to other lexicase approximations we also compare to plexicase selection [^1] Fixed a bug in the downsampling implementation in the released version of [10][10]. For plexicase selection, we use the hyperparameter setting $\alpha=1$. We use the Code-Building Genetic Programming (CBGP) system developed by Helmuth et al. [37] and test on a suite of problems from the PSB1 Benchmark [23] on which lexicase selection has shown good performance. We follow the default settings in CBGP, evolving 1000 individuals for 300 generations and using UMAD [20] with a rate of 0.09 as the sole variation operator. We present results in both the full data and downsampled paradigms. In the full data paradigm, all individuals are evaluated on all training cases. In the downsampled paradigm, a subset of the training cases is sampled at each generation and used to evaluate individuals in that generation. Downsampling has been shown to benefit the performance of Genetic Programming as we can increase the number of generations and/or individuals in the population for a given total computational budget [24, 2, 26]. Additional downsampling methods utilizing information from the population to choose the downsampled cases have been proposed [4], but we only study the randomly downsampled paradigm. In our experiments, we use a downsampling rate of 25%. The specific problems we study are Compare String Lengths (CSL), Median, Number IO, Replace Space with Newline (RSWN), Smallest, Vector Average, and Negative to Zero (NTZ).
Following the recommendation of the PSB1 benchmark, we report problem-solving performance as number of successful runs out of 100. We define a successful run as one which produces an individual with zero error on both the training cases and the unseen testing cases.
Table 1: Success rates of GP runs on six benchmark problems in the full data paradigm. The success rates across the board are very similar, and no results were significantly different between the selection methods. We perform chi-squared tests and underline results that were significantly worse than lexicase with ($p<0.05$). No results were significantly better than lexicase.
@@
CSL 0 0 0 0 0 0
@@
Table 1 compares the success rates of DALex, plexicase, and lexicase selection on six benchmark problems under the full and downsampled paradigms. No results were significantly different between the selection methods, and DALex achieves very similar success rates to lexicase selection. We find that DALex runs faster than lexicase selection but slower than plexicase selection. Detailed results can be found in the appendix.
To compare the selection dynamics of the three methods, we solve a GP problem using lexicase selection, and at each selection event we sample 50,000 individuals using each selection method. From this we build up the empirical probability distributions $\{p_{it}\}$ the probability of selecting individual $i$ at generation $t$ via lexicase selection, $\{p_{it}^{\prime}\}$ the probability of selecting individual $i$ at generation $t$ via DALex, and $\{p_{it}^{\prime\prime}\}$ the probability of selecting individual $i$ at generation $t$ via plexicase. To quantify the differences in probability distributions, we use the Jensen-Shannon divergence metric
@@
Where $q_{it}$ is $p^{\prime}_{it}$ or $p^{\prime\prime}_{it}$. Furthermore, for each run in which lexicase selection produces a generalizing individual, we track that individual’s lineage to find its ancestor $a_{t}$ in each generation $t$, of which there is only one per generation since we only use the single-parent variation operator UMAD. For each of these individuals we compute the probability ratio $r_{t}=\frac{q_{a_{t}t}}{p_{a_{t}t}}$ quantifying how much more likely DALex is to select the successful lineage than lexicase selection. We report the average of these two metrics over all generations and all runs.
@@
Figure 1: Fidelity of the DALex approximation to lexicase selection on CBGP problems. Error bars show the bootstrapped 95% confidence intervals. The probability ratio of DALex selecting the successful lineage is very close to 1 in all problems except for Vector Average. The Jensen-Shannon divergence of the DALex selection distribution from the lexicase selection distribution is close to 0 in all problems except for Vector Average and Negative To Zero.
@@
Figure 1 shows the results for these two metrics on the problems studied. In almost all problems, the probability of selecting the successful lineage is almost the same between DALex and lexicase. In contrast, the probability ratio of plexicase and lexicase deviates much more from 1 on many of the problems, indicating a less faithful approximation. Furthermore, the Jensen-Shannon divergences between the DALex and lexicase selection probability distributions are all very close to zero and much smaller than the divergences of the plexicase selection distribution from the lexicase selection distribution, indicating that DALex gives a close approximation of the lexicase selection dynamics. Finally, we expect increasing the value of the standard deviation to improve the fidelity of the DALex approximation.
@@
As an ablation experiment, we examine the effect of the distribution from which importance scores are chosen on the problem-solving ability of DALex. We compare our default selection from a normal distribution to selection from a uniform distribution and shuffling a range of evenly spaced numbers. In each case the importance scores are scaled so that they have a standard deviation of 20 prior to softmaxing.
@@
Table 2: Success rates of GP runs using different distributions for DALex. All distributions were scaled to a standard deviation of 20 unless otherwise noted. We perform chi-squared tests and underline results that were significantly worse than the normal distribution with ($p<0.05$). No results were significantly better than the normal distribution.
@@
4.2 Image Classification
To test the ability of DALex to mimic lexicase selection even with many training cases, we compare using DALex versus lexicase selection as the selection method in gradient lexicase. We use the popular VGG16 [39] and ResNet18 [16] neural networks on the CIFAR10 [30] dataset. This dataset is comprised of 32x32 bit RGB images distributed evenly across 10 classes. There are 50,000 training instances and 10,000 test instances. As shown in Ding et al. [9], lexicase selection often uses up to the entire training set to perform selection, especially towards the end of the genetic algorithm.
Gradient lexicase [11] is a hybrid of lexicase selection with deep learning, in which a population of deep neural networks (DNN) is evolved to fit an image classification dataset. In gradient lexicase, stochastic gradient descent on a subset of the training cases is used as the mutation operator. Lexicase selection proceeds by evaluating the population of DNNs on one training case at a time and keeping only the DNNs that correctly predict the training case’s label. In contrast, DALex evaluates each individual on the entire training set, assigns to each training case an error value of 0 if the case was correctly predicted or 1 otherwise, and computes a weighted sum of the error values. For this domain, we increase the particularity pressure tenfold to 200.
[…]
Table 3: Performance of five selection methods on the CIFAR10 dataset using the VGG16 architecture. DALex acheives similar test accuracy and lower runtime compared to lexicase selection.
Selector VGG16 ResNet18
Test Accuracy Runtime (h) Test Accuracy Runtime (h)
Vanilla SGD $92.6\pm 0.1$ $1.0\pm 0.001$ $93.72\pm 0.03$ $1.0\pm 0.006$
Lexicase $93.7\pm 0.05$ $13.1\pm 0.5$ $95.0\pm 0.1$ $35\pm 1$
DALex accuracy (std=200) $93.8\pm 0.1$ $8.27\pm 0.05$ $95.2\pm 0.1$ $9.0\pm 0.1$
DALex losses (std=200) $93.9\pm 0.1$ $8.25\pm 0.01$ $95.2\pm 0.2$ $8.9\pm 0.1$
We find that DALex performs just as well as lexicase selection whether selecting on cross-entropy losses or accuracy, and additionally runs much faster. This is likely due to lexicase selection’s evaluation of individuals on each training case with a batch size of 1. The small batch size causes too much communication overhead between the cpu and gpu, resulting in a slower runtime. Furthermore, as training progresses and the neural nets approach 100% accuracy on the training set, lexicase selection begins to exhibit its worst-case runtime. As the ResNet18 model reaches higher training set accuracy earlier than the VGG16 model, lexicase selection takes more than twice as long to run. With DALex, however, both models have a consistent runtime that is much lower than the runtime of lexicase selection.
4.3 SRBench
[…]
The statistics we examine in this paper, following the SRBench defaults, are test $R^{2}$, training time, and model complexity. We report the median of these statistics over all 20 problems, using bootstrapping to obtain a 95% confidence interval.
[…]
(b) Runtime
(c) Model Size
[…]
4.4 Learning Classifier Systems
Finally, to show that DALex can reproduce the success of batch-lexicase, we use the learning classifier system (LCS) from Aenugu et al. [3] We choose the led24 dataset [5] for the stark contrast between the performances of lexicase selection and batch-lexicase selection. The led24 problem [5], taken from the UCI ML Repository, tests the ability of selection methods to cope with noise. This dataset, consisting of 3200 instances, contains 7 boolean attributes corresponding to whether each of 7 LED lights in a display is lit up. It has 10 categories, the digits 0 through 9, corresponding to which digit is shown on the LED display. In this dataset, each boolean attribute has a 0.1 probability having its value inverted. The evolved rule distribution for this problem is sparse, with most rules representing fewer than 5 data instances [9]. This property makes the led24 problem unsuitable for lexicase selection and motivates the development of batch-lexicase selection. Since batched lexicase selection is also a relaxed version of lexicase selection, just like epsilon lexicase, for this domain we also need to determine the particularity pressure via a hyperparameter search, and choose the value of 20.
(a) Accuracy (Full)
(b) Runtime (Full)
(c) Rulecount (Full)
(d) Accuracy (Partial)
(e) Runtime (Partial)
(f) Rulecount (Partial)
Figure 3: Performance of four selection methods on the led24 problem. Bootstrapped 95% confidence intervals are indicated by error bars in the bar plots and shading in the line plots. DALex equals the highest test accuracy while having very low runtime and generating the fewest rules out of the four methods.
5 Conclusion and Future Work
While we have used reasonable presets for the standard deviation hyperparameter in DALex without performing significant hyperparameter search, we do not anticipate that a single hyperparameter setting will work for all cases or even for all generations during the course of a GP run. Therefore, future work could adaptively determine the optimal standard deviation during the course of evolution using hyperparameter search or even a more radical paradigm such as autoconstruction [41, 43], Additionally, since DALex computes a randomly weighted fitness function, it can be combined with selection methods other than pure elitist selection, such as tournament or fitness-proportional selection [13] to increase diversity. Furthermore, the weighted aggregation formulation of DALex is not confined to parent selection algorithms. This randomly weighted averaging concept could potentially be used to transform the losses for each training case in deep learning. By introducing more stochasticity into training, this could enable models to be trained with large batch sizes without the performance degradation that comes from lack of stochasticity [28]. Alternatively, DALex could be used to randomly weight each node in the policy head of a reinforcement learner when sampling trajectories, improving the exploration of off-policy learning algorithms such as Double Q-learning [15] or SAC [14] and potentially even removing the need for entropy regularization in policy-gradient algorithms.
Appendix 0.A Program Synthesis Runtime Comparisons
During our program synthesis experiments, we record the time taken to select the parents of the next generation. We then take the average of this metric over all generations and all runs and report them in Figure 4. DALex takes up to 5x less time to run than lexicase selection. While having a similar or slower runtime compared to plexicase selection, DALex has shown better performance as presented in Table 1, where plexicase selection occasionally produces significantly worse results than lexicase selection.

### 09 Feb 2024 v2

@@ 3.3 Modifications
For domains amenable to a relaxed lexicase variant such as epsilon-lexicase or batch-lexicase selection, we first standardize the population errors on each training case to have zero mean and unit variance. While not necessary, we find that this normalization step helps DALex perform well across many problems with a single hyperparameter setting.
Of the problem domains we study, Learning Classifier Systems (LCS) [46] has a significantly different structure and therefore requires further modification. Since an LCS individual represents a rule that matches a subset of the training cases, each individual will only have errors defined on a subset of the training cases. For these problems, we let the individual have error 0 on cases for which it is undefined. We also use the individual’s support vector, which has a 1 for each training case on which the individual is defined, and a 0 otherwise. To compute the individual’s average error, we take the dot product of the error vector with the weight vector and then normalize by dividing this value by the dot product of the support vector with the weight vector.
@@
  • $\bullet$
    $E=\{e_{i,j}\}$ the error value of individual $i$ on training case $j$, or 0 if individual $i$ is not defined on training case $j$
  • $\bullet$
    $N=\{s_{i,j}\}$ the support of individual $i$ on training case $j$, which equals 1 if individual $i$ is defined on case $j$, otherwise 0
  • $\bullet$
    $n$ the number of selection events, $m$ the number of training cases
  • $\bullet$
    particularity_pressure, the standard deviation of the sampled weights
  • $\bullet$
    Relaxed, whether to simulate a relaxed version of lexicase selection.
Result:
  • $\bullet$
    ${\tt idx}$, the $n$ indices of the selected individuals to be the parents of the next generation.
if Relaxed then
@@
Algorithm 1 Diversely Aggregated Lexicase Selection, batched selection event
In problems where individuals are defined on all training cases, the support is 1 everywhere, so the normalization constant $SW^{T}$ is simply the sum of the training case weights, which is 1 due to the softmax operation. Under this condition, our algorithm degenerates to the simpler case considered above. The complete algorithm pseudocode, with the augmentations described above to accommodate for LCS and SR problems, is described in Algorithm 1. In our pseudocode, we define arithmetic operations as the vectorized element-wise operations used in libraries like numpy, and batch together the $n$ importance score vectors for the $n$ individual selection events into an $n\times m$ matrix.
@@ 4.1 CBGP
We first compare DALex to lexicase selection on program synthesis problems, the class of problem for which lexicase selection was first developed. While we use a particularity pressure of 20 for these experiments to illustrate the robustness of DALex, we recommend setting the particularity pressure as large as possible within the range of floating point precision, or around 200, for problems suitable for lexicase selection. As DALex can exactly simulate lexicase selection in the limit of infinite particularity pressure, we also assess the ability of DALex to replicate the lexicase selection probability distribution with a moderate particularity pressure. To situate our results with respect to other lexicase approximations we also compare to plexicase selection [^1] Fixed a bug in the downsampling implementation in the released version of [8][8]. For plexicase selection, we use the hyperparameter setting $\alpha=1$. We use the Code-Building Genetic Programming (CBGP) system developed by Helmuth et al. [37] and test on a suite of problems from the PSB1 Benchmark [21] on which lexicase selection has shown good performance. For an overview on CBGP, see appendix A. We follow the default settings in CBGP, evolving 1000 individuals for 300 generations and using UMAD [18] with a rate of 0.09 as the sole variation operator. We present results in both the full data and downsampled paradigms. In the full data paradigm, all individuals are evaluated on all training cases. In the downsampled paradigm, a subset of the training cases is sampled at each generation and used to evaluate individuals in that generation. Downsampling has been shown to benefit the performance of Genetic Programming as we can increase the number of generations and/or individuals in the population for a given total computational budget [23, 22, 25]. Additional downsampling methods utilizing information from the population to choose the downsampled cases have been proposed [2], but we only study the randomly downsampled paradigm. In our experiments, we use a downsampling rate of 25%. The specific problems we study are Compare String Lengths (CSL), Median, Number IO, Replace Space with Newline (RSWN), Smallest, Vector Average, and Negative to Zero (NTZ).
Following the recommendation of the PSB1 benchmark, we report problem-solving performance as number of successful runs out of 100. We define a successful run as one which produces an individual with zero error on both the training cases and the unseen testing cases.  [^2] Due to specific quirks of the code-building system, it is very difficult for CBGP to generalize successfully on Compare String Lengths.
Table 1: Success rates of GP runs on six benchmark problems in the full data and downsampled paradigms. The success rates across the board are very similar, and no results were significantly different between the selection methods. We perform chi-squared tests following [8] and underline results that were significantly worse than lexicase selection with ($p<0.05$). No results were significantly better than lexicase selection.
@@
CSL [^2] Due to specific quirks, Compare String Lengths is systemically difficult for CBGP to solve. 0 0 0 0 0 0
@@
Table 1 compares the success rates of DALex, plexicase, and lexicase selection on six benchmark problems under the full and downsampled paradigms. No results were significantly different between the selection methods, and DALex achieves very similar success rates to lexicase selection. For our ablation experiments on the distribution of the sampled importance scores, see appendix C. Additionally, we find that DALex runs faster than lexicase selection but slower than plexicase selection. Detailed results can be found in appendix B.
To compare the selection dynamics of the three methods, we solve the same PSB1 problems using lexicase selection, and at every generation we sample 50,000 individuals using each selection method. From this we build up the empirical probability distributions $\{p_{it}\}$ the probability of selecting individual $i$ at generation $t$ via lexicase selection, $\{p_{it}^{\prime}\}$ the probability of selecting individual $i$ at generation $t$ via DALex, and $\{p_{it}^{\prime\prime}\}$ the probability of selecting individual $i$ at generation $t$ via plexicase selection. To quantify the differences in probability distributions between lexicase selection and its approximations, we use the Jensen-Shannon divergence metric
@@
Where $q_{it}$ is $p^{\prime}_{it}$ or $p^{\prime\prime}_{it}$ for DALex and plexicase selection, respectively. The lower the JS-divergence of the DALex/plexicase selection probability distribution from the lexicase selection probability distribution, the better the approximation to lexicase selection. Furthermore, for each run in which lexicase selection produces a generalizing individual, we track that individual’s lineage to find its ancestor $a_{t}$ in each generation $t$, of which there is only one per generation since we only use mutation operators. For each of these individuals we compute the probability ratio $r_{t}=\frac{q_{a_{t}t}}{p_{a_{t}t}}$ quantifying how much more likely DALex and plexicase selection are to select the successful lineage than lexicase selection. We report the average of these two metrics over all generations and all runs.
@@
Figure 1: Fidelity of the DALex and plexicase approximations to lexicase selection on CBGP problems. Error bars show the bootstrapped 95% confidence intervals. The ratio of selecting the successful lineage via DALex versus the probability using lexicase is very close to 1 in 6 out of 7 problems. The Jensen-Shannon divergence of the DALex selection probability distribution from the lexicase distribution is close to 0 in 5 out of 7 problems.
@@
Figure 1 shows the results for these two metrics on the problems studied. In almost all problems, the probability of selecting the successful lineage is almost the same between DALex and lexicase selection. In contrast, the probability ratio of plexicase and lexicase selection often deviates much more from 1, indicating a less faithful approximation. Furthermore, the Jensen-Shannon divergences between the DALex and lexicase selection probability distributions are often very close to zero and much smaller than the corresponding divergences of plexicase selection from lexicase selection, indicating that DALex gives a close approximation of lexicase selection dynamics. As mentioned before, we expect increasing the value of the standard deviation to improve the fidelity of the DALex approximation, bringing the DALex JS divergences even lower and the DALex probability ratios even closer to 1.
4.2 Image Classification
To test the ability of DALex to mimic lexicase selection even with many training cases, we compare DALex against lexicase selection as the selection method in gradient lexicase selection. We use the popular VGG16 [39] and ResNet18 [14] neural networks on the CIFAR10 [29] dataset. This dataset is comprised of 32x32 bit RGB images distributed evenly across 10 classes. There are 50,000 training instances and 10,000 test instances. As shown in Ding et al. [7], lexicase selection often uses up to the entire training set to perform selection, especially towards the end of the genetic algorithm.
Gradient lexicase selection[9] is a hybrid of lexicase selection with deep learning, in which a population of deep neural networks (DNN) is evolved to fit an image classification dataset. In gradient lexicase selection, stochastic gradient descent on a subset of the training cases is used as the mutation operator. Lexicase selection proceeds by evaluating the population of DNNs on one training case at a time and keeping only the DNNs that correctly predict the training case’s label. In contrast, DALex evaluates each individual on the entire training set, assigns to each training case an error value of 0 if the case was correctly predicted or 1 otherwise, and computes a weighted sum of the error values. As this domain is intended to test the limits of our lexicase approximation with its large number of training cases, we use the previously recommended particularity pressure of 200.
[…]
Table 2: Performance of three selection methods and two architectures on the CIFAR10 dataset. Vanilla SGD is also included as a baseline. DALex acheives similar test accuracy and lower runtime compared to lexicase selection.
Selector VGG16 ResNet18
Test Accuracy Runtime (h) Test Accuracy Runtime (h)
Vanilla SGD $92.8\pm 0.2$ $0.962\pm 0.003$ $93.8\pm 0.1$ $1.003\pm 0.006$
Lexicase $93.7\pm 0.1$ $11.1\pm 0.3$ $95.1\pm 0.2$ $29\pm 1$
DALex accuracy (std=200) $93.7\pm 0.2$ $8.7\pm 0.1$ $95.3\pm 0.1$ $9.1\pm 0.1$
DALex losses (std=200) $93.7\pm 0.2$ $8.66\pm 0.02$ $95.2\pm 0.2$ $9.2\pm 0.1$
We find that DALex performs just as well as lexicase selection whether selecting on cross-entropy losses or accuracy, and additionally runs much faster. This is likely due to the fact that lexicase selection evaluates individuals on each training case with a batch size of 1. The small batch size causes too much communication overhead between the cpu and gpu, resulting in a slower runtime. Furthermore, as training progresses and the models approach 100% accuracy on the training set, lexicase selection begins to exhibit its worst-case runtime. As the ResNet18 model reaches a higher training set accuracy than the VGG16 model, lexicase selection takes more than twice as long to run. With DALex, however, both models have consistent and much lower runtimes.
4.3 SRBench
[…]
The statistics we examine in this paper, following the SRBench defaults, are test $R^{2}$, training time, and model complexity. Unlike in program synthesis, we find that the runtime of epsilon lexicase dominates the algorithm runtime, so we expect the choice of selection method to significantly impact the overall runtime. We report the median of these statistics over all 20 problems, using bootstrapping to obtain a 95% confidence interval.
[…]
(b) Runtime
(c) Model Size
[…]
4.4 Learning Classifier Systems
Finally, to show that DALex can reproduce the success of batch-lexicase, we use the learning classifier system (LCS) from Aenugu et al. [1] We choose the led24 dataset [3] for the stark contrast between the performances of lexicase selection and batch-lexicase selection. The led24 problem [3], taken from the UCI ML Repository, tests the ability of selection methods to cope with noise. This dataset, consisting of 3200 instances, contains 7 boolean attributes corresponding to whether each of 7 LED lights in a display is lit up. It has 10 categories, the digits 0 through 9, corresponding to which digit is shown on the LED display. In this dataset, each boolean attribute has a 0.1 probability having its value inverted. The evolved rule distribution for this problem is sparse, with most rules representing fewer than 5 data instances [7]. This property makes the led24 problem unsuitable for lexicase selection and motivates the development of batch-lexicase selection. For this domain, we use a particularity pressure of 20 as determined by a preliminary hyperparameter search.
(a) Accuracy (Full)
(b) Runtime (Full)
(c) Rulecount (Full)
(d) Accuracy (Partial)
(e) Runtime (Partial)
(f) Rulecount (Partial)
Figure 3: Performance of four selection methods on the led24 problem in the full data and partial data scenarios. Bootstrapped 95% confidence intervals are displayed as error bars for the accuracy and runtime plots and as a shaded region for the rulecount plot. DALex has approximately equal test accuracy to batch lexicase while having much lower runtime and generating the fewest rules.
5 Conclusion and Future Work
While we have used reasonable presets for the standard deviation hyperparameter in DALex, we do not anticipate that a single hyperparameter setting will work for all cases or even for all generations during the course of a GP run. Therefore, future work could determine the optimal standard deviation using hyperparameter search or even a more radical paradigm such as autoconstruction [41, 43], Additionally, since DALex computes a randomly weighted fitness function, it can be combined with selection methods other than pure elitist selection, such as tournament or fitness-proportional selection [11] to increase diversity. In fact, the weighted aggregation formulation of DALex is not confined to parent selection algorithms. The concept of randomly weighted aggregation could also be used to transform the losses for each training case in deep learning. By introducing more stochasticity into training, this could enable models to be trained with large batch sizes without suffering performance degradation [27]. Alternatively, DALex could be used to randomly weight each node in the policy head of a reinforcement learner when sampling trajectories, improving the exploration of off-policy learning algorithms such as Double Q-learning [13] or SAC [12].
6 Acknowledgements
This work was performed in part using high-performance computing equipment at Amherst College obtained under National Science Foundation Grant No. 2117377. Any opinions, findings, and conclusions or recommendations expressed in this publication are those of the authors and do not necessarily reflect the views of the National Science Foundation. The authors would like to thank Ryan Boldi, Bill Tozier, Tom Helmuth, Edward Pantridge and other members of the PUSH lab for their insightful comments and suggestions.
Appendix 0.A Code-Building GP for Program Synthesis
Code Building Genetic Programming (CBGP)[37] is a typed genetic programming system that aims to use constructs like those found in human programming to solve program synthesis problems. Incorporating concepts like data types, data structures, and control flow into a GP system not only promises to create more complex hierarchical systems but also produces more interpretable programs which are analogous in structure to human created programs.
CBGP encodes programs as directed acyclic graphs (DAGs) where each leaf node is an input or a constant, and each inner node a function. Together, these nodes are called “expressions" and form a computational graph. In these graphs, the return type of child nodes are required to be sub-types of the corresponding argument in the parent node for type safety. The CBGP implementation used in this paper supports the expression types Constants, Inputs, Functions, Methods, Constructors, and Higher Order Functions.
While CBGP programs are represented by a graph structure, it still uses the linear plushy genome representation found in other GP systems like PushGP. This representation consists of a sequence of expression tokens representing program syntax and structure tokens representing the hierarchical structure of the program. These plushy genomes are compiled into Push programs which are then translated into CBGP DAGs for execution. As in PushGP, CBGP uses the Uniform Mutation by Addition and Deletion (UMAD) mutation operator[18] as the sole variation operator.
Appendix 0.B Program Synthesis Runtime Comparisons
During our program synthesis experiments, we record the time taken to select the parents of the next generation. We then take the average of this metric over all generations and all runs and report them in Figure 4. The results show that DALex is up to 5x faster than lexicase selection. Despite having a similar or slower runtime compared to plexicase selection, DALex is a more faithful approximation to lexicase selection and has shown better performance as presented in Table 1, where plexicase selection sometimes produces worse results than lexicase selection.
Appendix 0.C Ablation of DALex probability distributions
@@
As an ablation experiment, we examine the effect of the distribution from which importance scores are chosen on the problem-solving ability of DALex. We compare our default selection from a normal distribution to selection from a uniform distribution and shuffling a range of evenly spaced numbers. The range case is interesting in that it is guaranteed to produce lexicase selection when the spacing between weights is larger than the magnitude difference between the largest possible error and the smallest possible nonzero error. In each case the importance scores are scaled so that they have a standard deviation of 20 prior to softmaxing.
@@
Table 3: Success rates of GP runs using different distributions for DALex. All distributions were scaled to a standard deviation of 20 unless otherwise noted. We perform chi-squared tests following [8] and underline results that were significantly worse than the normal distribution with ($p<0.05$). No results were significantly better than the normal distribution.